OpenStudy (anonymous):

find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (psymon):

Is that meant to be cos^2(x) or actually cos2x?

OpenStudy (anonymous):

cos ^2x

OpenStudy (anonymous):

if you substitute u for cos x what is the equation?

OpenStudy (psymon):

Ah, okay. Well, start off by pretending that those aren't cosines, but actually x's and factor like a quadratic.

OpenStudy (anonymous):

Replace cos(x) with u. Solve for u

OpenStudy (anonymous):

So it would be u^2 +2u+1=0

OpenStudy (psymon):

Yep.

OpenStudy (anonymous):

right now how does that factor?

OpenStudy (anonymous):

u ( u+2) +1 = 0

OpenStudy (anonymous):

would that be it

OpenStudy (anonymous):

no...you can't factor only part of a polynomial. Trinomials usually factor to two binomials. Do you know how to factor trinomials?

OpenStudy (anonymous):

I dont know i remembering it just forgot it

OpenStudy (anonymous):

remember it

OpenStudy (anonymous):

so x^2 + 2x + 1 ... I usually consider what could possibly multiply to give me x^2 (the first term) and then what could multiply to give me +1 (the last term) then I test all the possible combinations. In this case it is pretty easy because there is only one combinAtion.

OpenStudy (anonymous):

x*x = x^2 and +1*+1=+1 or -1*-1 = +1. So it is either (x+1) (x+1) or (x-1)(x-1).

OpenStudy (anonymous):

But (x-1)(x-1) = x^2-2x+1 and (x+1) (x+1) = x^2+2x+1 so it is (x+1) (x+1)

OpenStudy (anonymous):

so since u^2+2u+1=(u+1)(u+1)=0 , you know either u+1 =0 or u+1 =0 which is redundant.

OpenStudy (anonymous):

Sine u+1=0 then substitute cos x back in for u and solve from there.

OpenStudy (anonymous):

I would strongly suggest reviewing factoring when you are studying this part of trigonometry. It will make your life much easier.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Could you still help me on trying trying to find all solutions to it ?

OpenStudy (anonymous):

Ok so you have that (u+1)(u+1)=0 ...so we are really looking for U+1=0.

OpenStudy (anonymous):

SO this means that u= -1 but remember u = cos x so cos x = -1. Now based on what you know from the unit circle when does cos x = -1?

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

yep but you have to consider all solutions so ever revolution that would land at that same spot so... \[x=\pi + 2\pi n \] where \[n \in \mathbb{Z}\] the line above is simply saying that n is an integer so that only when the value is pi or any revolution that lands on pi (i.e. all co terminal angles to pi)